# Probing Environmental Dependence of High-Redshift Galaxy Properties with the Marked Correlation Function

### v1

@@ 5 Summary and Conclusions
  • The MCF of LBGs using the UV magnitude and color show stronger deviation from unity compared to $r-$ band luminosity derived MCFs of $z\sim 0$ galaxy samples from SDSS and GAMA (Skibba et al., 2013; Sureshkumar et al., 2021).
@@
Despite similar absolute magnitudes, galaxy samples exhibit significant variations in dark matter halo and galaxy properties - such as large-scale bias and effective halo mass- across redshifts. These variations could explain the observed differences in marked clustering signals. To control for these effects, we compared MCFs of samples at different redshifts with comparable effective large-scale galaxy bias and average halo mass (defined in Sections 4.4 and 4.5), leading to the below findings:
@@ Acknowledgements
We are grateful to the anonymous referee for the detailed and thoughtful comments. We thank Yuichi Harikane for providing us with galaxy samples from PDR2 for analysis and clarifying doubts. CJ is supported by the University Grant Commission, India, through a BSR start-up grant (F.30-463/2019(BSR)) and the Rashtriya Uchchatar Shiksha Abhiyan scheme (RUSA 2.0, project number: T3A). EM acknowledges the financial support in the form of the fellowship of Cochin University of Science and Technology, Kerala and RUSA 2.0 (project number: T3A).

### v2

@@ 5 Summary and Conclusions
  • The MCFs of LBGs using the UV magnitude and color show a stronger deviation from unity compared to the $r-$ band luminosity derived MCFs of $z\sim 0$ galaxy samples from SDSS and GAMA (Skibba et al., 2013; Sureshkumar et al., 2021).
@@
Despite similar absolute magnitudes, galaxy samples exhibit significant variations in dark matter halo and galaxy properties - such as large-scale bias and effective halo mass- across redshifts. These variations could explain the observed differences in marked clustering signals. To control for these effects, we compared the MCFs of samples at different redshifts with comparable effective large-scale galaxy bias and average halo mass (defined in Sections 4.4 and 4.5), leading to the following findings.
@@ Acknowledgements
We are grateful to the anonymous referee for the detailed and thoughtful comments. The authors thank Yuichi Harikane for providing us with galaxy samples from PDR2 for analysis and clarifying doubts. CJ is supported by the University Grant Commission, India, through a BSR start-up grant (F.30-463/2019(BSR)) and the Rashtriya Uchchatar Shiksha Abhiyan scheme (RUSA 2.0, project number: T3A). EM acknowledges the financial support in the form of the fellowship of Cochin University of Science and Technology, Kerala and RUSA 2.0 (project number: T3A).